# Antidistillation Sampling

### v2

@@ Abstract
Frontier models that generate extended reasoning traces inadvertently produce rich token sequences that can facilitate model distillation. Recognizing this vulnerability, model owners may seek sampling strategies that limit the effectiveness of distillation without compromising model performance. Antidistillation sampling provides exactly this capability. By strategically modifying a model’s next-token probability distribution, antidistillation sampling poisons reasoning traces, rendering them significantly less effective for distillation while preserving the model’s practical utility.
@@ 1 Introduction
Large language models (LLMs) trained to produce extended reasoning traces demonstrate impressive performance across math, coding, and general reasoning benchmarks [e.g 1]. These generated traces, however, serve a dual purpose—not only do they enhance model capabilities, but they also facilitate model distillation, wherein a secondary model learns to replicate the original model’s abilities by training on its generated traces [2, 3]. Notably, model distillation enables substantial capability gains in the secondary model at a fraction of the computational cost required for training a similarly performant model from scratch [3].
Despite the benefits of model distillation, the effectiveness and efficiency of this technique present several downsides for companies serving frontier reasoning models. First, returning extended reasoning traces constitutes a forfeiture of valuable intellectual property, enabling competitors to cheaply replicate frontier capabilities. Second, the possibility of distillation incentivizes frontier model providers to hide token probabilities, summarize reasoning traces, or, more generally, restrict user-model interaction. And finally, safe model behaviors (e.g., resistance to jailbreaking attempts [4, 5]) are often not inherited by distilled models, enabling the generation of objectionable content [6].
@@
To address these issues, we introduce antidistillation sampling—a method designed to reduce the effectiveness of model distillation. The core idea underpinning antidistillation sampling is to adjust a reasoning model’s sampling distribution so that generated traces (1) poison distillation attempts while (2) maintain high likelihood under the original, unadjusted distribution. This approach protects proprietary capabilities while preserving the original model’s utility for downstream applications.
Figure 1: Reasoning traces generated via antidistillation sampling poison distillation attempts, while simultaneously preserving the teacher’s downstream performance. The top and bottom rows shows results for MATH [7] and GSM8K [8], respectively. The left columns shows teacher accuracies under different sampling schemes, and the right column shows the performance of students distilled on traces sampled via these different sampling schemes. Notably, for the same teacher performance on both datasets, antidistillation sampling significantly degrades the performance of the distilled models relative to temperature sampling.
1.1 Related Work
Various frontier AI labs acknowledge and benefit from the effectiveness of model distillation. For instance, OpenAI offers model distillation as a service within their API [9]. The recognition of model distillation’s potential, which motivates such pipelines, dates back at least to Schmidhuber [10]. More recently,  Hinton et al. [2] demonstrated that distilled “specialist” models can achieve impressive performance in multiple domains. Since then, a growing body of work aims to understand the transference of capabilities via distillation [11, 12, 13, 14]. In fact, there is speculation that some labs train commercial-facing LLMs in part via distillation, possibly by harvesting extended reasoning traces generated from models belonging to their direct competitors [15]. This practice constitutes a strategic vulnerability for frontier model maintainers, given the demonstrated value of these reasoning traces, and underscores the importance of the algorithm we introduce in this work.
@@ 1.1 Related Work
The threat model we address in this paper—where a student model is trained on data generated by a teacher model—intersects with several aspects of model privacy and security. For instance, model extraction attacks acquire weights via query-level access without training or distillation [16] and training data extraction attacks harvest training data directly from frontier models [17]. While antidistillation sampling may offer some protection against these attacks, such analysis remains beyond our current scope. More relevant is the literature on data poisoning, where maliciously crafted data is injected into a model’s training set to induce specific unintended downstream effects (see, e.g., [18]). Rando and Tramèr [19] even show the effectiveness of adding backdoors to preference data, sabotaging models fine-tuned with RLHF. Our contribution bridges data poisoning and privacy techniques to protect the valuable knowledge encoded in frontier models.
@@
Finally, we position antidistillation sampling within the broader framework of controlled decoding for LLMs [20], where supplementary objectives steer the decoding process. Existing approaches in this domain include leveraging contrastive objectives to enhance generation quality [21], reformulating constrained decoding as an optimization problem [22], and incorporating energy-based constraints [23]. While related, antidistillation sampling solves a different problem: by implementing a new, distillation-aware penalization term in the decoding objective, our approach poisons generated reasoning traces to undermine the performance of models fine-tuned on these outputs.
@@ 2 Antidistillation Sampling
To introduce and derive antidistillation sampling, we first provide a high-level sketch, followed by the desiderata outlining the desired qualities for poisoning distillation attempts. We then derive the antidistillation sampling method in §2.2. Algorithm 1 summarizes the key steps involved in implementing our approach.
An overview of our approach.
The core objective of antidistillation sampling is to adjust a model’s next-token distribution to balance two competing goals: sampling tokens with high likelihood under the original, unadjusted distribution and sampling tokens that effectively poison distillation attempts. Throughout, we refer to the model from which reasoning traces are sampled as the teacher, and the model being distilled as the student. We begin by introducing notation describing sampling from the teacher’s unadjusted distribution. Next, we quantify how model distillation impacts the student model’s performance on a given downstream task. This analysis yields a key insight—we can incorporate this performance metric directly into the teacher’s sampling distribution. This takes the form of a directional derivative capturing the change in the teacher’s sampling distribution along the update direction in the student’s weight space. However, due to the high cost needed to compute this directional derivative, the final portion of our derivation identifies an efficient finite-difference approximation for this term, which is inexpensive to compute and, as we demonstrate in §3, results in effective distillation poisoning.
2.1 Preliminaries
To formalize our presentation of antidistillation, we first introduce notation. We consider an LLM to be a mapping from a sequence of input tokens $x_{1:t}=(x_{1},\dots,x_{t})$ to a distribution over the next token, where each token $x_{j}$ is an element of a vocabulary set $\mathcal{V}=\{1,\dots,V\}$. This distribution is parameterized by weights $\theta$ and can be expressed as follows.
$p(\cdot|x_{1:t};\theta):\mathcal{V}^{t}\to[0,1]^{V}$ (1)
We write $p(\cdot|x_{1:t};\theta)$ to denote the distribution of all next-token probabilities, whereas $p(x_{t+1}|x_{1:t};\theta)$ refers to the scalar probability of a given next token $x_{t+1}$. Typically, tokens are generated according to a scaled version of this distribution by sampling as follows: [^1] Variants of this sampling scheme include top-$k$ sampling (i.e., limiting sampling to the tokens with the top-$k$ largest probabilities), greedy sampling (i.e., sampling from the same objective while letting $\tau\rightarrow 0$), and beam search, but we focus mainly on temperature-based sampling here.
Desiderata for antidistillation sampling.
Model distillation involves a student language model—parameterized by $\theta_{S}$, with a distribution over next tokens given by $p(\>\cdot\>|x_{1:t};\theta_{S})$—trained on data generated from a teacher model parameterized by $\theta_{T}$. These models do not need to share the same parameter space, and therefore the parameter vectors $\theta_{S}$ and $\theta_{T}$ need not be comparable; indeed, in many cases, a student model may have substantially fewer parameters than the teacher.
The aim of antidistillation sampling is to generate tokens from the teacher $\theta_{T}$, which perform well according to a metric used to evaluate teacher samples, while simultaneously having the property that training on these tokens cannot improve performance on this same task. In more detail, we aim to adjust the teacher’s sampling procedure to achieve the following objectives simultaneously:
  1. I.
  2. II.
Proxy models.
2.2 Deriving Antidistillation Sampling
$=\theta_{P}+\eta\nabla_{\theta_{P}}\log p(x_{t+1}|x_{1:t};\theta_{P})$ (4)
If $\Delta(x_{t+1}|x_{1:t})$ is positive, the update in eq. 4 increases the loss; if $\Delta(x_{t+1}|x_{1:t})$ is negative, the update decreases the loss. Thus, our goal is to adjust the teacher’s sampling distribution so that tokens sampled from the teacher both have (1) high likelihood under the teacher’s unadjusted distribution and (2) yield larger (i.e., more positive) values of $\Delta$.
where $Z$ is a normalization term appropriately scaled (relative to eq. 2) to accommodate the penalty, and $\lambda>0$ is a regularization coefficient that facilitates a trade-off between sampling from the teacher’s distribution and sampling tokens that maximally increase student’s downstream loss. Unfortunately, directly implementing eq. 6 is impractical, as we would need to compute $\Delta(x_{t+1}|x_{1:t})$ for each potential next token $x_{t+1}\in\mathcal{V}$, requiring $V$ gradients to be computed as well as $V$ evaluations of the downstream loss $\ell$, which, in turn, is assumed to involve a lengthy computation to produce.
Figure 2: Antidistillation sampling uses a tunable parameter $\lambda$ to control the trade-off between teacher accuracy and distillability. The baseline involves sampling from the teacher with increasing temperature $\tau$ to show that we can produce traces that are bad for distillation at some cost in teacher accuracy. One important feature of the blue temperature sampling curve is that to bring the student accuracy down below the undistilled accuracy, the teacher performance has to drop to 20%. On the other hand, with antidistillation sampling, the teacher model can still get 70% accuracy while producing traces that bring the student’s performs down below the undistilled accuracy.
An efficient implementation of antidistillation sampling.
The core of our proposed approach is an efficient mechanism to approximate the sampling process above. As a starting point, observe that $\Delta(x_{t+1}|x_{1:t})$ can be scaled by a factor of $1/\eta$ without changing the relative penalties for each $x_{t+1}$ (i.e., we could fold this term into the $\lambda$ regularization penalty). Then, by taking the limit of $\Delta(x_{t+1}|x_{1:t})/\eta$ as $\eta\to 0$, we have that
That is, the limit is the inner product between the gradient $\nabla_{\theta_{P}}\log p(x_{t+1}|x_{1:t};\theta_{P})$ and the downstream loss gradient $\nabla\ell(\theta_{P})$. Notice that the expression in eq. 8 no longer involves the evaluation of the downstream loss for each token in $\mathcal{V}$. Rather, $\nabla\ell(\theta_{P})$ can be computed and stored once, after which the only remaining task is to efficiently evaluate eq. 8 for each token $x_{t+1}\in\mathcal{V}$. To do so, the key observation is that the directional derivative is symmetrical. Thus, we can rewrite eq. 8 as a finite difference limit in the other term, i.e., in terms of a finite difference update to $\log p(x_{t+1}|x_{1:t};\theta_{P})$. This gives
$=\lim_{\epsilon\rightarrow 0}\frac{\log p(x_{t+1}|x_{1:t};\theta_{P}+\epsilon\nabla\ell(\theta_{P}))-\log p(x_{t+1}|x_{1:t};\theta_{P})}{\epsilon}$ (10)
$\widehat{\Delta}(\>\cdot\>|x_{1:t})$ $=\frac{\log p(\>\cdot\>|x_{1:t};\theta_{P}+\epsilon\nabla\ell(\theta_{P}))-\log p(\>\cdot\>|x_{1:t};\theta_{P})}{\epsilon}$ (11)
for some appropriately chosen small value of $\epsilon$, [^3] It is also possible to compute the inner product of this gradient exactly via an explicit Jacobian-vector product (not to be confused with a vector-Jacobian product used in backpropagation). However, even in modern automatic differentiation frameworks, Jacobian-vector products tend lack support for a handful of important operations, such as SDPA. Meanwhile flash attention [24] only supports Float16 and BFloat16. Thus, we use finite differences for the improved convenience and efficiency, but we verify the correctness of the finite difference approximation against autograd in §C.1. where $\widehat{\Delta}(x_{t+1}|x_{1:t})$ approaches eq. 8 for all next tokens $x_{t+1}$ in the limit as $\epsilon\to 0$. Finally, we sample according to the teacher’s adjusted sampling distribution:
@@ 2.2 Deriving Antidistillation Sampling
Input: Prompt $x_{1:n}$, max tokens $N$, penalty multiplier $\lambda$, approximation parameter $\epsilon$, temperature $\tau$
@@ 2. For each token index $t=n,n+1,\ldots,N-1$:
3 Empirical Results
Architectures and benchmarks.

### v3

@@ Abstract
Frontier models that generate extended reasoning traces inadvertently produce token sequences that can facilitate model distillation. Recognizing this vulnerability, model owners may seek sampling strategies that limit the effectiveness of distillation without compromising model performance. Antidistillation sampling provides exactly this capability. By strategically modifying a model’s next-token probability distribution, antidistillation sampling poisons reasoning traces, rendering them significantly less effective for distillation while preserving the model’s utility. Our code is available at https://github.com/locuslab/antidistillation-sampling.
@@ 1 Introduction
Large language models (LLMs) trained to produce reasoning traces have achieved strong performance on math, coding, and reasoning benchmarks Jaech et al. (2024); Anthropic (2024); Guo et al. (2025). These traces, however, serve a dual purpose. They not only enhance model performance, but also enable distillation, a process by which a secondary model replicates the original model’s capabilities by training on its generations Hinton et al. (2015); Schmidhuber (1992); Gou et al. (2021). Notably, distillation can result in substantial capability gains at a fraction of the computational cost needed to train similarly performant models from scratch.
While effective and efficient, the viability of distillation poses several downsides for companies deploying frontier reasoning models. First, returning reasoning traces represents a forfeiture of intellectual property, which can allow competitors to cheaply replicate frontier capabilities. Second, the threat of distillation incentivizes limiting user access by obscuring token probabilities or truncating reasoning traces. Finally, model safety is often not preserved by distillation, which enables the generation of harmful content Sabbaghi et al. (2025); Burgess and Newman .
@@
To address these issues, we introduce antidistillation sampling (see Figure˜1). The main idea underpinning antidistillation sampling is to adjust a model’s sampling distribution so that generated traces maintain high likelihood under the unadjusted distribution, and distillation attempts are simultaneously poisoned. To operationalize this idea, we first formulate the general problem of poisoning reasoning models trained via distillation. We then derive one solution to this problem (see Algorithm 1), which facilitates a precise trade-off between two competing objectives—the utility of the original model and the effectiveness of distillation poisoning—while incurring minimal computational overhead.
To illustrate our empirical results, consider a reasoning model that achieves 72% accuracy on MMLU. Naively distilling this model using greedy sampling can produce a student that reaches up to 52% accuracy. If the teacher model increases its sampling temperature, its accuracy slightly decreases (e.g., by 4%), yet the distilled student’s accuracy remains largely unchanged at 52%. In contrast, using antidistillation sampling with the same 4% reduction in teacher accuracy significantly reduces the student’s capabilities, lowering its accuracy to as low as 40%. Our findings on GSM8K Cobbe et al. (2021), MATH Hendrycks et al. (2021), and MMLU Hendrycks et al. (2020) indicate that model owners can effectively limit distillation quality via antidistillation sampling.
Figure 1: Reasoning traces generated via antidistillation sampling poison distillation attempts while simultaneously preserving the teacher’s performance. The teacher’s logits are perturbed in a direction $\Delta$, leading to samples that significantly degrade distilled model performance relative to naive temperature sampling. For more details, see Figure˜2 and §3.
2 Related work
Model distillation. The prominence and effectiveness of model distillation—and, more generally, model compression Buciluǎ et al. (2006); Cheng et al. (2017); Polino et al. (2018)—is rooted in a rich literature dating back to Schmidhuber (1992), and, more recently, to Hinton et al. (2015). Since these seminal works, a growing body of literature has sought to benchmark the performance of distilled models and to algorithmically maximize the effectiveness of distillation Shridhar et al. (2023); Li and Mooney (2024); Gu et al. (2023a); Goldblum et al. (2020); Romero et al. (2014). Indeed, a variety of frontier AI labs have incorporated distillation as a core technique, both to efficiently enable frontier capabilities Times (2025); DeepSeek-AI (2025) and to improve model safety via context distillation Bai et al. (2022); Snell et al. (2022); Askell et al. (2021). However, this practice constitutes a strategic vulnerability for frontier model maintainers, given the demonstrated value of these reasoning traces (see, e.g., Jaech et al. (2024); Guan et al. (2024); Lightman et al. (2023); Uesato et al. (2022)).
@@ 2 Related work
Model security. The setting we address in this paper—where a student model is trained on data generated by a teacher model—intersects with several aspects of model security. For instance, model extraction attacks acquire weights via query-level access, whereas training data extraction attacks are designed to harvest training data Carlini et al. (2024); Nasr et al. (2023). While antidistillation sampling may offer some protection against these attacks, such analysis remains beyond our scope. More relevant is the literature on data poisoning, where maliciously crafted data is injected into a model’s training set to induce specific downstream effects (see, e.g., Goldblum et al. (2022)). In this vein, Rando and Tramèr (2024) show the effectiveness of adding backdoors to preference data, sabotaging LLMs finetuned with RLHF. Our contribution bridges data poisoning and privacy techniques to protect the valuable knowledge encoded in frontier models.
Distillation prevention. A related line of work has sought to develop algorithms that prevent distillation. In the context of computer vision, Ma et al. (2021) corrupt the teacher’s logits via self-training, whereas follow-up work shows that returning only the top-$k$ logits tends to harm distillation Ma et al. (2022). Also related are watermarking algorithms, which seek to adjust model logits to detect whether a model has been distilled Zhao et al. (2023); Gu et al. (2023b); Sander et al. (2024); Kirchenbauer et al. (2023). And while this family of methods preserves teacher accuracy, their static nature—each input generally yields a deterministic logit vector—presents security vulnerabilities: a distiller can learn an inverse transformation by saving input-output pairs, and thereby fine-tune to recover uncorrupted logits. To mitigate this shortcoming, Chen et al. (2025) propose a session-dynamic defense that monitors the sensitivity of a user’s queries and perturbs the logits once a threshold is crossed. In contrast, antidistillation sampling is fully dynamic; it perturbs each token’s distribution on-the-fly using gradients from a hidden proxy model, turning generation into a moving target in a similar fashion to cryptographic stream ciphers.
@@
Language model decoding. Finally, we position antidistillation sampling within the broader framework of controlled decoding for LLMs Mudgal et al. (2023), where supplementary objectives steer the decoding process. Existing approaches in this domain include using contrastive objectives to enhance generation quality Li et al. (2022), reformulating constrained decoding as an optimization problem Ji et al. (2023), and incorporating energy-based constraints Qin et al. (2022). While related, antidistillation sampling solves a different problem: by implementing a new, distillation-aware penalization term in the decoding objective, our approach poisons generated reasoning traces to undermine the performance of models fine-tuned on these outputs.
@@ 3 Antidistillation sampling
Input: Prompt $x_{1:n}$, max tokens NN, penalty multiplier $\lambda$, approximation parameter $\epsilon$, temperature $\tau$
@@ 2. For each token index $t=n,n+1,\ldots,N-1$:
To motivate antidistillation sampling, we first sketch a high-level overview of our problem setting in §3.1. Based on this setting, we provide a desiderata outlining the desired qualities for poisoning distillation attempts in §3.2. We then derive the antidistillation sampling method (summarized in Algorithm 1) in §3.3.
3.1 An overview of antidistillation
The core objective of antidistillation sampling is to adjust a model’s next-token distribution to balance two competing goals: sampling tokens with high likelihood under the original, unadjusted distribution and sampling tokens that effectively poison distillation attempts. Throughout, we refer to the model from which reasoning traces are sampled as the teacher, and the model being distilled as the student.
Our derivation relies on quantifying how model distillation impacts the student model’s performance on a given downstream task. This analysis yields a key insight—we can incorporate this performance metric directly into the teacher’s sampling distribution. This takes the form of a directional derivative capturing the change in the teacher’s sampling distribution along the update direction in the student’s weight space. However, due to the high cost needed to compute this directional derivative, the final portion of our derivation identifies an efficient finite-difference approximation for this term, which is inexpensive to compute and, as we demonstrate in §4, results in effective distillation poisoning.
3.2 Preliminaries
We consider an LLM to be a mapping from a sequence of input tokens $x_{1:t}=(x_{1},\dots,x_{t})$ to a distribution over the next token, where each token is an element of a vocabulary set $\mathcal{V}=\{1,\dots,V\}$. This distribution is parameterized by weights $\theta$ and can be expressed as $p(\cdot|x_{1:t};\theta)$. We write $p(\cdot|x_{1:t};\theta)$ to denote the distribution of all next-token probabilities, whereas $p(x_{t+1}|x_{1:t};\theta)$ refers to the scalar probability of a given next token $x_{t+1}$. Typically, tokens are generated according to a scaled version of this distribution: [^1] Variants of this sampling scheme include top-$k$ sampling (i.e., limiting sampling to the tokens with the top-$k$ largest probabilities), greedy sampling (i.e., sampling from the same objective while letting $\tau\rightarrow 0$), and beam search, but we focus mainly on temperature-based sampling here.
Desiderata for antidistillation. Distillation involves a student model—parameterized by $\theta_{S}$, with a distribution over next tokens given by $p(\>\cdot\>|x_{1:t};\theta_{S})$—trained on data generated from a teacher model parameterized by $\theta_{T}$. These models do not need to share the same parameter space, and therefore the parameter vectors $\theta_{S}$ and $\theta_{T}$ need not be comparable; indeed, a student model may have substantially fewer parameters than the teacher.
The aim of antidistillation sampling is to generate tokens that perform well according to a metric used to the evaluate teacher, while simultaneously having the property that training on these tokens cannot improve performance on this same task. In more detail, we aim to adjust the teacher’s sampling procedure to simultaneously satisfy the following:
  1. I.
  2. II.
Proxy models.
Figure 2: An illustration of approximating $\Delta$. The teacher model performs antidistillation sampling autoregressively, based on its perturbed distribution by $\Delta$. Given an input prompt and $t$ reasoning tokens from the teacher, $\Delta$ is approximated by the difference of the log probability of each token in the vocabulary between two copies of the proxy model (created by performing a single gradient ascent step using the downstream task loss on the proxy model); this difference is represented by the   area in the bar plot.
3.3 Deriving antidistillation sampling
$=\theta_{P}+\eta\nabla_{\theta_{P}}\log p(x_{t+1}|x_{1:t};\theta_{P})$ (3)
If $\Delta(x_{t+1}|x_{1:t})$ is positive, the update in eq.˜3 increases the loss; if $\Delta(x_{t+1}|x_{1:t})$ is negative, the update decreases the loss. Thus, our goal is to adjust the teacher’s sampling distribution so that tokens sampled from the teacher both have (1) high likelihood under the teacher’s unadjusted distribution and (2) yield larger (i.e., more positive) values of $\Delta$.
where $Z$ is a normalization term appropriately scaled (relative to eq.˜1) to accommodate the penalty, and $\lambda>0$ is a regularization coefficient that facilitates a trade-off between sampling from the teacher’s distribution and sampling tokens that maximally increase student’s downstream loss. Unfortunately, directly implementing eq.˜5 is impractical, as we would need to compute $\Delta(x_{t+1}|x_{1:t})$ for each potential next token $x_{t+1}\in\mathcal{V}$, requiring $V$ gradients to be computed as well as $V$ evaluations of the downstream loss $\ell$, which, in turn, is assumed to involve a lengthy computation to produce.
An efficient implementation. The core of our proposed approach is an efficient mechanism to approximate the sampling process above. As a starting point, observe that $\Delta(x_{t+1}|x_{1:t})$ can be scaled by a factor of $1/\eta$ without changing the relative penalties for each $x_{t+1}$ (i.e., we could fold this term into the $\lambda$ regularization penalty). Then, by taking the limit of $\Delta(x_{t+1}|x_{1:t})/\eta$ as $\eta\to 0$, we have that
That is, the limit is the inner product between the gradient $\nabla_{\theta_{P}}\log p(x_{t+1}|x_{1:t};\theta_{P})$ and the downstream loss gradient $\nabla\ell(\theta_{P})$. Notice that the expression in eq.˜7 no longer involves the evaluation of the downstream loss for each token in $\mathcal{V}$. Rather, $\nabla\ell(\theta_{P})$ can be computed and stored once, after which the only remaining task is to efficiently evaluate eq.˜7 for each token $x_{t+1}\in\mathcal{V}$. To do so, the key observation is that the directional derivative is symmetrical. Thus, we can rewrite eq.˜7 as a finite difference limit in the other term, i.e., in terms of a finite difference update to $\log p(x_{t+1}|x_{1:t};\theta_{P})$. This gives
$\qquad\qquad=\lim_{\epsilon\rightarrow 0}\frac{\log p(x_{t+1}|x_{1:t};\theta_{P}+\epsilon\nabla\ell(\theta_{P}))-\log p(x_{t+1}|x_{1:t};\theta_{P}-\epsilon\nabla\ell(\theta_{P}))}{2\epsilon}$ (9)
$\widehat{\Delta}(\>\cdot\>|x_{1:t})$ $=\frac{\log p(\>\cdot\>|x_{1:t};\theta_{P}+\epsilon\nabla\ell(\theta_{P}))-\log p(\>\cdot\>|x_{1:t};\theta_{P}-\epsilon\nabla\ell(\theta_{P}))}{2\epsilon}$ (10)
for some appropriately chosen small value of $\epsilon$, where $\widehat{\Delta}(x_{t+1}|x_{1:t})$ approaches eq.˜7 for all next tokens $x_{t+1}$ in the limit as $\epsilon\to 0$. Intuitively, $\widehat{\Delta}(x_{t+1}|x_{1:t})$ measures how much sampling token $x_{t+1}$ would degrade a proxy student’s performance after a single update. Finally, we sample according to the teacher’s adjusted sampling distribution:
4 Exploring antidistillation in practice
Figure 3: Antidistillation sampling uses a tunable parameter $\lambda$ to control the trade-off between teacher accuracy and distillability. The baseline involves sampling from the teacher with increasing temperature $\tau$ to show that we can produce traces that are bad for distillation at some cost in teacher accuracy. One important feature of the blue temperature sampling curve is that to bring the student accuracy down below the undistilled accuracy, the teacher performance has to drop to 20%. On the other hand, with antidistillation sampling, the teacher model can still get 70% accuracy while producing traces that bring the student’s performance down below the undistilled accuracy.
Through a range of experiments, we demonstrate the effectiveness of antidistillation sampling and discuss several interesting phenomena. First, we show that the hyperparameter $\lambda$ provides model owners with precise control over the trade-off between nominal utility and non-distillability. This trade-off persists across various teacher-student model configurations, and notably, remains effective even when the proxy student is from a different model family than the actual student—validating the practical applicability of our method in realistic scenarios where the model owners lack knowledge of potential student architectures. Additionally, we address methodological questions through a pointed empirical analysis.
4.1 Experimental setup
First, we detail our selection of model architectures and benchmark datasets, chosen to represent realistic distillation scenarios across varied reasoning tasks. Next, we describe the computation of $\nabla\ell(\theta_{P})$ (eq.˜10). Finally, we outline our baseline comparison methodology.